Two squares have a combined area of 68cm squared. the side of one square is for times as long as a side of the other square. Find the perimeter of each square
Answer by ewatrrr(24785)

 
Hi,
.
A = Lw
s^2 + (4s)(4s) = 68cm^2
17s^2 = 68cm^2
s^2 = 4 (Tossing out the negative solution for unit measure)
s = 2
P = 8cm and P = 32cm